# How to Create a Flat Ten or Eleven Dimensional Space-time in the Interior of an Asymptotically Flat Four Dimensional String Theory

### v2

@@ 5 Discussions
In [1] we had described a general strategy for producing almost locally flat background where the moduli fields take a different value compared to the values they take at infinity. In this paper we have described a particular example of this mechanism where starting from a four dimensional string theory with finite values of all the moduli fields, we have described a background that has a large locally flat region where the space-time is effectively ten dimensional and the string coupling remains finite and another large locally flat region where the space-time is effectively eleven dimensional.
@@
This of course does not demonstrate that we can reach all possible values of the moduli this way. It will be interesting to try to establish this or find the limitations if any. In particular the moduli space of string theories have interesting topology changing transitions and one would like to realize them as part of space-time background. One interesting case of these is the conifold transition[7, 8] – that takes a deformed conifold to a resolved conifold. One could ask if it is possible to construct a background where in some region of space-time the internal manifold is a resolved conifold and in another region it is a deformed conifold. Since this involves both vector multiplet and hypermultiplet moduli of $N=2$ supersymmetric string theory, the configuration is likely to involve both black holes and loops of black strings.
@@
Finally we note that while the asymptotic geometry is taken to be flat space-time (times an internal manifold) the geometries that we can construct in the interior are not limited to flat space-time. For example is ten dimensional type IIB string theory we can take a stack of $N$ spherical D3-branes so that it has finite energy, but take the radius of the sphere to be very large so that the relaxation time is large. In that case the near horizon geometry of a local patch of the D3-branes will be $AdS_{5}\times S^{5}$ with five form flux[9] and we can study properties of this background by sending observers close to the horizon. [^6] If we scale the 4-form field strength according to (2.5) then the three-brane charge will grow as $\lambda^{4}$[1] and we shall get a locally flat space-time. To get $AdS_{5}\times S^{5}$ geometry we need to keep the three brane charge fixed. Of course if we try to approach infinitely close to the horizon any experiment will take infinite amount of time from the asymptotic observer’s point of view due to the red shift, but once we fix the accuracy with which we want to do the experiment and estimate the time that will be needed from the asymptotic observer’s point of view to perform such an experiment, we can take the size of the 3-brane system to be sufficiently large so that it does not evolve appreciably during the period of the experiment. Similarly by taking large stacks of spherical M2-branes or M5-branes in M-theory we can produce $AdS_{4}\times S^{7}$ or $AdS_{7}\times S^{4}$ geometries[9]. Since both, the ten dimensional type IIB string theory and the eleven dimensional M-theory, can be produced as backgrounds in the interior of an asymptotically flat four dimensional string theory, obtained by compactifying type IIA string theory on $T^{6}$, we see that all the $AdS_{p}\times S^{q}$ backgrounds listed above can be considered as different states of the same underlying theory.

### v3

@@ 6 Discussions
In [1] we had described a general strategy for producing almost locally flat background where the moduli fields take a different value compared to the values they take at infinity. In this paper we have described a particular example of this mechanism where starting from a four dimensional string theory with finite values of all the moduli fields, we have described backgrounds that have large locally flat regions where the space-time is effectively ten dimensional and the string coupling remains finite and another background that has a large locally flat region where the space-time is effectively eleven dimensional.
@@
This of course does not demonstrate that we can reach all possible values of the moduli this way. It will be interesting to try to establish this or find the limitations if any. In particular the moduli space of string theory has interesting topology changing transitions and one would like to realize them as part of space-time background. One interesting case of these is the conifold transition[7, 8] – that takes a deformed conifold to a resolved conifold. One could ask if it is possible to construct a background where in some region of space-time the internal manifold is a resolved conifold and in another region it is a deformed conifold. Since this involves both vector multiplet and hypermultiplet moduli of $N=2$ supersymmetric string theory, the configuration is likely to involve both black holes and loops of black strings.
@@
Finally we note that while the asymptotic geometry is taken to be flat space-time (times an internal manifold) the geometries that we can construct in the interior are not limited to flat space-time. For example in ten dimensional type IIB string theory we can take a stack of $N$ spherical D3-branes so that it has finite energy, but take the radius of the sphere to be very large so that the relaxation time is large. In that case the near horizon geometry of a local patch of the D3-branes will be $AdS_{5}\times S^{5}$ with five form flux[9] and we can study properties of this background by sending observers close to the horizon. [^5] If we scale the 4-form field strength according to (2.5) then the three-brane charge will grow as $\lambda^{4}$[1] and we shall get a locally flat space-time. To get $AdS_{5}\times S^{5}$ geometry we need to keep the three brane charge fixed. Of course if we try to approach infinitely close to the horizon then any experiment will take infinite amount of time from the asymptotic observer’s point of view due to the red shift, but once we fix the accuracy with which we want to do the experiment and estimate the time that will be needed from the asymptotic observer’s point of view to perform such an experiment, we can take the size of the 3-brane system to be sufficiently large so that it does not evolve appreciably during the period of the experiment. Similarly by taking large stacks of spherical M2-branes or M5-branes in M-theory we can produce $AdS_{4}\times S^{7}$ or $AdS_{7}\times S^{4}$ geometries[9]. Since both, the ten dimensional type IIB string theory and the eleven dimensional M-theory, can be produced as backgrounds in the interior of an asymptotically flat four dimensional string theory, obtained by compactifying type IIA string theory on $T^{6}$, we see that all the $AdS_{p}\times S^{q}$ backgrounds listed above can be considered as different states of the same underlying theory. In fact, in this case, we could also wrap the branes along the compact cycles instead of taking them to be spherical since from the asymptotic observer’s viewpoint these are particle like states carrying large but finite energy and charges.
@@
Appendix A Electrically charged black hole solution in type II on $T^{6}$
In this appendix we shall use the results of [6] to describe the non-rotating black hole solutions in type II string theory on $T^{6}$ carrying fundamental string winding and momentum charges. [^6] Even though the solution in [6] was given for heterotic string theory, the result in type II string theory can be obtained easily from there by truncating the 28 dimensional charge vector to 12 dimensional charge vector and the $28\times 28$ matrix valued scalar moduli $M$ to a $12\times 12$ symmetric $SO(6,6)$ matrix. The solution given here is obtained from the one in [6] by setting to zero the rotation parameter $a$. Also $\Phi$ in [6] is our $2\phi_{4}$. The four dimensional string metric $ds_{4}^{2}$, the four dimensional dilaton $\phi_{4}$ the gauge fields $A_{\mu}$, and the moduli of $T^{6}$, described by a $12\times 12$ symmetric $O(6,6)$ matrix $M$, take the form:
$e^{-2\phi_{4}}$ $=$ $e^{-2\phi_{0}}\,\Delta^{1/2}\,\rho^{-2}\,,$
$M$ $=$ $I_{12}+\pmatrix{Pnn^{T}&Qnp^{T}\cr Qpn^{T}&Ppp^{T}}\,,$
$A_{t}$ $=$ $-{1\over\sqrt{2}}\,m\,\rho\,\Delta^{-1}\,\pmatrix{\sinh\alpha\{\cosh\beta\rho^{2}+m\rho(\cosh\alpha-\cosh\beta)\}\vec{n}\cr\sinh\beta\{\cosh\alpha\rho^{2}+m\rho(\cosh\beta-\cosh\alpha)\}\vec{p}}\,,$ (A.1)
where
$\displaystyle\Delta$$\equiv$ $\rho^{4}+2m\rho^{3}(\cosh\alpha\cosh\beta-1)+m^{2}\rho^{2}(\cosh\alpha-\cosh\beta)^{2}\,,$
$P$$\equiv$ $2\,\Delta^{-1}\,m^{2}\rho^{2}\sinh^{2}\alpha\sinh^{2}\beta\,,$
$Q$$\equiv$ $-2\Delta^{-1}m\rho\sinh\alpha\sinh\beta\{\rho^{2}+m\rho(\cosh\alpha\cosh\beta-1)\}\,.$ (A.2)
[…]
${1\over 2}\pmatrix{R^{2}+R^{-2}&R^{2}-R^{-2}\cr R^{2}-R^{-2}&R^{2}+R^{-2}}\,.$ ) . (A.3)